A plane moving at a velocity of 800mm travels in7200min. Find the planes acceleration in mm/mins


1
Expert's answer
2020-10-23T11:52:53-0400

"a=\frac{2s}{t^2}\\a=\frac{2(800)}{7200^2}=0.00003\frac{mm}{min^2}"